OpenStudy (anonymous):

Integration using tables of integrals... Integrate e^2x arctan(e^x) dx

OpenStudy (anonymous):

I said u = e^x so du would be e^x as well but im not sure if this is the right substitution...

OpenStudy (helder_edwin):

\[ \large \int e^{2x}\arctan(e^x)\,dx \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

do u know integration by parts?

OpenStudy (anonymous):

yes... Int udv = uv -int vdu

OpenStudy (helder_edwin):

ok

OpenStudy (anonymous):

you have to use int by parts? it says using table tho...

OpenStudy (helder_edwin):

i have nothing like that in my table.

OpenStudy (anonymous):

what about \[\int\limits u \tan ^{-1} u du\]

OpenStudy (helder_edwin):

if u have it. then u should use it.

OpenStudy (anonymous):

i just got stumpd because if i let the u = e^x then i cant sub it in for the e^2x, you know?

OpenStudy (helder_edwin):

i know. we could use this \[ \large u=\arctan e^x \] then \[ \large \tan u=e^x \]

OpenStudy (helder_edwin):

give me few minutes. i'll work this with pencil on paper

OpenStudy (anonymous):

ok ill attempt as well

OpenStudy (anonymous):

I have the answer, the textbook ...

OpenStudy (anonymous):

It says you should get \[\frac{ (e^{2x} +1) }{ 2 } \arctan(e ^{x}) - \frac{ e^x }{ 2 } +c \]

OpenStudy (anonymous):

this is what you get if you use that formula i had above..... now, in order top get this answer the u would have to be e^x

OpenStudy (helder_edwin):

this is what i got.

OpenStudy (helder_edwin):

let \(u=\arctan e^x\) so \(\tan u=e^x\). from this we have that \(\tan^2u=(e^x)^2=e^{2x}\) also \(\ln\tan u=x\) and from this \[ \large dx=\frac{1}{\tan u}\sec^2u\,du \] ok so far?

OpenStudy (anonymous):

ok..

OpenStudy (helder_edwin):

now \[ \large \int e^{2x}\arctan e^x\,dx= \int \tan^2u\cdot u\cdot\frac{\sec^2u}{\tan u}\,du \] \[ \large =\int\tan u\cdot u\cdot\frac{1}{\cos^2u}\,du =\int\frac{\sin u}{\cos u}\cdot u\cdot\frac{1}{\cos^2u}\,du \] \[ \large =\int\frac{u\cdot\sin u}{\cos^3u},du \] ok?

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

now using integration by parts \[ \large \begin{alignat*}{2} A&=u &\qquad dB&=\frac{\sin u}{\cos^3u}\,du\\ Da&=du &\qquad B&=\frac{1}{2\cos^2u} \end{alignat*} \]

OpenStudy (helder_edwin):

\[ \large \int\frac{u\cdot\sin u}{\cos^3u}\,du= \frac{u}{2\cos^2u}-\int\frac{1}{2\cos^2u}\,du \] \[ \large =\frac{1}{2}\left[u\cdot\sec^2u-\int\sec^2u\,du\right] \] \[ \large =\frac{1}{2}\left[u\cdot\sec^2u-\tan u\right] \] OK?

OpenStudy (anonymous):

ok...

OpenStudy (helder_edwin):

now replacing \(u=\arctan e^x\) we have finally \[ \large \int e^{2x}\arctan e^x\,dx=\frac{1}{2} \left[\arctan e^x\cdot\sec^2(\arctan e^x)- \tan(\arctan e^x)\right] \]

OpenStudy (helder_edwin):

but \[ \large \tan(\arctan e^x)=e^x \] and using \(\sec^2z=\tan^2z+1\) we have \[ \large \sec^2(\arctan e^x)=\tan^2(\arctan e^x)+1=(e^x)^2+1= e^{2x}+1 \]

OpenStudy (helder_edwin):

therefore \[ \large \int e^{2x}\arctan e^x\,dx=1/2[(e^{2x}+1)\cdot\arctan e^x-e^x] \]

OpenStudy (anonymous):

ahhhhhh

OpenStudy (helder_edwin):

it is weird that they asked u to do this "by tables".

OpenStudy (helder_edwin):

see u some other time.

OpenStudy (anonymous):

thank you very much

OpenStudy (helder_edwin):

u r welcome